40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 25 19:42:12 2014, ericq, Update, General, Vent progress 
    Reply  Fri Sep 26 12:25:34 2014, ericq, Update, General, Vent progress 
       Reply  Fri Sep 26 16:10:14 2014, ericq, Update, General, Vent update 
          Reply  Fri Sep 26 17:13:39 2014, ericq, Update, General, Vent update 
             Reply  Mon Sep 29 12:47:51 2014, ericq, Update, General, Vent update 
                Reply  Mon Sep 29 17:10:51 2014, ericq, Update, General, Vent update 
                   Reply  Mon Sep 29 18:12:24 2014, ericq, Update, General, Vent update 
                      Reply  Tue Sep 30 11:53:29 2014, ericq, Update, General, Vent update 
                         Reply  Tue Sep 30 16:18:57 2014, Steve, Update, VAC, pump down #78 has started pd78hastarted.png
                            Reply  Tue Sep 30 18:02:53 2014, Steve, Update, VAC, pump down #78 stops at 320 Torr 3h.png
                               Reply  Wed Oct 1 10:21:19 2014, Steve, Update, VAC, pump down #78 stops at 3.5 Torr pd78secondStop.png
                                  Reply  Wed Oct 1 16:26:53 2014, Steve, Update, VAC, pump down #78 completed pd78completedin25hrs.png
                                     Reply  Thu Oct 2 09:23:23 2014, Steve, Update, VAC, vent 77 vent77SRMmoved.png
                                        Reply  Fri Oct 3 10:10:37 2014, Steve, Update, VAC, cold cathode gauge reading switched CC1switched.png
                                           Reply  Fri May 8 14:30:09 2015, Steve, Update, VAC, CC1 cold cathode gauges are baked now cc1bake95C.jpg
                                              Reply  Tue May 12 14:57:52 2015, Steve, Update, VAC, CC1 cold cathode gauges are baked now 
Message ID: 10545     Entry time: Fri Sep 26 16:10:14 2014     In reply to: 10544     Reply to this: 10546
Author: ericq 
Type: Update 
Category: General 
Subject: Vent update 

Today so far:

  • I moved SRM forward by 3mm
  • Then I leveled the ITMY table 
  • At this point, bringing the ITMY oplev beam back onto its QPD got me back to green locking and IR flashes 
  • AS and POY beams are both making it out to their tables, as seen by IR card. (Though not to their in-air optics)

Here's my quick brain dump of things to do before we can pump down (anyone see anything missing?):

  • Check the clearance of the POY beam at the SRM cage
  • Re-do distance reconstruction measurements, confirm desired SRC length
  • Lock the SRM cage down fully (right now, has 2 clamps on, and one laying unused)
  • Align SRM for SRC flashes
  • Adjust SRM OSEM positions as needed
  • Adjust SRM oplev beam path, measure lever arm for calibration
  • Confirm beam spots on output mirrors in ITMY and BS chambers are ok
  • Take pictures of ITMY chamber. 
  • Closeup checklist
ELOG V3.1.3-